OpenStudy (anonymous):

y=2x²-3 y=3x-1 HOW do you find the solutions?

OpenStudy (anonymous):

@AravindG

OpenStudy (johnweldon1993):

see how y = two different things? set them equal to each other

OpenStudy (anonymous):

huh?

OpenStudy (aravindg):

both are equivalent to y so equate them !!

OpenStudy (johnweldon1993):

y=2x²-3 y=3x-1 see how y = ...2 different things? well by logic...that must mean that those 2 things are equal as well correct? so 2x² - 3 = 3x - 1 now solve for x

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

2x² - 3 = 3x - 1 +3 +3 2x²=3x+2 ÷2 ÷2 Would that be 1.5x+1? doesn't sound right...

OpenStudy (johnweldon1993):

2x² - 3 = 3x - 1 add 1 to both sides 2x² - 2 = 3x subtract 3x from both sides 2x² -3x - 2 = 0 look like anything familiar?

OpenStudy (anonymous):

yes it does. Thank you so much!

OpenStudy (johnweldon1993):

let us know what you get for your answers

OpenStudy (anonymous):

ok i think im going to finish tomorrow. I need to set up my birthday party really earily it its almost 1.... Will you be on tomorrow?

OpenStudy (johnweldon1993):

haha ....probably not tomorrow...but anyone else would be glad to help you! :)

OpenStudy (johnweldon1993):

and happy birthday :)

OpenStudy (anonymous):

Yeah i guess. Thanks!